Saharaqueensasa:

Adult tickets to an amusement park cost dollars. Children's tickets cost dollars. The Henson family bought 3 adult and 1 child tickets for $156. The Garcia family bought 2 adult and 3 child tickets for $153. Part 1 out of 2 Drag and drop each term into the correct box to complete the equations to represent the Hensons' cost and the Garcias' cost.

loganisbored:

?

loganisbored:

it doesnt say how much

Saharaqueensasa:

what do u mean

loganisbored:

nvm

loganisbored:

i cant help you on this i thought i could but ask @tetsxpreme

Saharaqueensasa:

ok thx any way

Florisalreadytaken:

"Adult tickets to an amusement park cost dollars. Children's tickets cost dollars. " since tose amounts of money are empty, im assuming thats what you have to find -- so off we start The Henson family bought `3 adult` and `1 child` tickets for `$156`. The Garcia family bought `2 adult` and `3 child` tickets for `$153`. let the adults tickets be \(x\), and the child tickets be \(y\) from that, we can write this system: \[ \begin{cases} 3x +y = $156 \\ 2x+3y= $153 \end{cases} \] do you know how to solve this system?

Saharaqueensasa:

Jed is graphing the design for a kite on a coordinate grid. The four vertices of the kite are at (− 4 3 , 2 3 ), ( 14 3 , − 4 3 ), ( 14 3 , − 16 3 ), ( 2 3 , − 16 3 ). One kite strut will connect points and . The other will connect points and . Find the point where the struts cross. Enter your answer using improper fractions in simplest form. The struts cross at .